The base of an isosceles triangle is 29 cm and the perimeter is 75 cm. How long is a leg?
Answer by jim_thompson5910(35256)

x+x+29 = 75

2x+29 = 75

2x = 75-29

2x = 46

x = 46/2

x = 23

Each congruent leg is 23 cm.